# Do Code LLMs Do Static Analysis?

### v1

@@ 3.7 Research Methods – RQ3
Our research method for answering RQ3 is to finetune the open-source LLMs and evaluate those models on static analysis tasks. We show those configurations in Tables 1 and 2. For example, we show that we finetune jam to do ast generation and evaluate jam with the same task. We do not finetune jam for callgraph generation because of the context length limit (see Section 3.4). Note that we increase the number of tokens to 10k for callgraph generation because callgraph generation would require larger context size. We use the same hyperparamters and metrics for evaluation as in RQ1 and RQ2.
@@ 3.8 Research Methods – RQ4
Our research method for answering RQ4 is to finetune the the open-source LLMs on downstream tasks after those models have been finetuned on static analysis tasks. We show those configurations in Tables 1 and 2. For example, in the first row of jam column in RQ4, we show that we first finetune jam on the AST generation task. Then, we further finetuned it for code summarization. We used the same hyperparameters and metrics described in RQ1.
@@ 3.9 Threats to Validity
We organize threats to validity into threats to internal validity, external validity, and construct validity based on Wohlin et al. (2012). The threats to internal validity are that different hyperparamters can have different results. We mitigate this threat by following the accompanying papers. The threats to construct validity are that our experimental results may not be able to generalize to different tasks. We mitigate this by using three different code development tasks and static analysis tasks alone with two different programming languages. The threats to external validity are that different models might generate different results. We mitigate this threat by using two different commercial LLMs and two different open-souce LLMs.
Table 4: Meteor and USE score for code summarization
Metrics
model config METEOR USE p_m p_u
Java jam base 40.95 68.80 - -
srcml 41.28 69.01 0.02 0.06
dataflow graph 40.09 68.32 0.99 0.99
callgraph - - - -
codellama base 51.44 74.56 - -
srcml 51.19 74.48 0.99 0.88
dataflow graph 51.27 74.54 0.94 0.62
callgraph 51.43 74.57 0.53 0.43
C codellama base 29.65 50.62 - -
srcml 29.96 50.85 0.99 0.99
dataflow graph 29.87 50.73 0.99 0.93
callgraph 29.97 50.91 0.99 0.99
Table 5: BLEU score for code generation
model config BLEU p
Java jam base 23.51 -
srcml 24.00 $p<0.01$
dataflow graph 22.79 0.63
callgraph - -
codellama base 31.42 -
srcml 30.40 0.92
dataflow graph 34.22 $p<0.001$
callgraph 28.93 0.99
C codellama base 48.20 -
srcml 48.18 0.70
dataflow graph 47.54 0.50
callgraph 47.82 0.73
Table 6: BLEU score for code translation
model config BLEU p
Java jam base 50.34 -
srcml 49.26 0.90
dataflow graph 50.25 0.08
callgraph - -
codellama base 80.36 -
srcml 80.52 0.56
dataflow graph 79.82 0.57
callgraph 81.84 $p<0.01$
C codellama base 37.82 -
srcml 37.50 0.84
dataflow graph 41.64 0.91
callgraph 39.70 0.91
@@ 4 Experimental Results
4.1 RQ1: Open Models, Code Dev. Tasks
We showed the results for RQ1 in Table 45, and  6. In Java, we found that jam has 40.95 METEOR score and 68.80 USE score for code summarization. Larger models with larger training data further improve the results across three different datasets. For example, we found that CodeLlaMA has 25% improvement in METEOR score and 8% improvement over USE in code summarization. To our surprise, code translation has over 50% improvement when we finetuned it with 13B parameter models. Although the result shows that larger LLMs tend to have the better results compared with smaller models (Zhang et al., 2024a), we cannot attribute the improvement on CodeLlaMA with the size of the model because of the data leakage issues in the training set of those large language models (López et al., 2024). However, we have confidence that those methods are not in the  jam model’s training set.
For code summarization and code translation, we observed the higher performance in Java and low performance in C/C++. The rationale for the higher performance on Java may be because of higher amount of training data as a result of the popularity of the programming language (Jiménez, 2024). Another possible explanation for higher performance on the code translation task in Java is that the syntax of Java and C# are similar to each other and models have more information from source language. However, C and Rust have larger difference in terms of syntax (Shiraishi and Shinagawa, 2024). For example, to define an integer $a=0$ in Java and C#, we can simply write “int a = 0;” in both languages. However, for C and Rust, we would write “int a = 0;” and “let a = 5;”. Overall, these findings showed that LLMs can perform better on tasks with similar syntax and more popular programming languages but perform worse on tasks with different syntax.
In terms of code generation, we observed the low performance in Java and higher performance in C/C++. The possible explanation is that LLMs struggles with code generation when it comes the code with more complicated logic (Zhong and Wang, 2024; Qiu et al., 2024). This result also aligns with recent findings that LLMs usually struggle with generating code on object-oriented programming languages (Jiménez, 2024; Cao et al., 2024). The cause of this difference may be because the inability of LLMs to infer the information from longer context although the context windows is long, e.g. 128k tokens for GPT or Gemini (Li et al., 2024a).
@@
dataflow graph gpt-base 0.001 0.004 0
gpt-in-context 0.011 0.029 1.29E-06
gemini-base 0.025 0.007 0
gemini-in-context 0.010 0.036 2.33E-04
codellama-finetuned 0.623 0.771 1.24E-01
jam-finetuned 0.550 0.780 2.50E-01
@@
dataflow graph gpt-base 0.0009 0.003 1.00E-04
gpt-in-context 0.004 0.010 4.00E-03
gemini-base 0.0016 0.0035 1.35E-03
gemini-in-context 0.007 0.017 6.00E-03
codellama-finetuned 0.120 0.217 8.00E-02
@@
4.2 RQ2: Closed Models, Static Analysis Tasks
We showed the results for AST generation for both Java and C/C++ in Table 8. In Java, we found that gpt-base has the Levenshtein score of 0.45 and gemini-base has the Levenshtein score of 0.8. We also found the similar results in gemini although we found higher results in gemini, which has Levenshtein score of 0.64 for gemini-base and Levenshtein score of 0.93 for gemini-in-context. For C/C++, we observed the similar trend as in Java that in-context learning helps to improve the overall results on AST generation. However, we observed the smaller improvement in C/C++ dataset. For example, we found around 80% improvement between GPT-base and GPT-in-context in Java dataset. However, we only found around 30% improvement between GPT-base and GPT-in-context in C/C++ dataset, which has around 50% difference in improvement. This result aligns with the finding that LLMs can have better performance on more popular programming languages (Jiménez, 2024). The gap between C/C++ and Java may imply the lack of reasoning skills on LLMs to generalize from one programming language to another programming language.
@@ 4.2 RQ2: Closed Models, Static Analysis Tasks
We showed the results for callgraph and data flow graph in Table 9 for both Java and C/C++. For callgraph generation in Java, we found that gpt-base has Jaccard similarity of 0.121 and gpt-in-context has Jaccard similarity of 0.122. gemini-base and gemini-in-context have slightly higher performance, which is 0.186 and 0.191 respectively. We also found that chain accuracy is the lowest among all other metrics. In C/C++, we found the similar trend that in-context learning does not help to improve the results and chain accuracy is the lowest among all other metrics. For data flow generation, we found the similar results on both Java and C/C++ dataset. Overall, we found that LLMs do a poor jobs on callgraph generation and data flow graph generation and in-context learning does not help to improve the performance for callgraph generation and data flow generation. The possible explanation is that LLMs do not have the fundamental ability for callgraph generation and data flow generation. It is also possible that more samples would help LLMs to learn more useful information for these two tasks (Xu and Zhang, 2024). Another possibility is that simply throwing everything into one giant prompt would not help to improve the model for more complex tasks (Zamfirescu-Pereira et al., 2023) although LLMs can take very long context.
Overall, we found that closed-source LLMs do not do static analysis very well. To our surprise, we found that in-context learning only helps to improve AST generation but not callgraph generation and data flow generation. The possible explanation is that it is easy for LLMs to learn the tags for AST generation. However, LLMs struggle with looking at larger context for callgraph generation or more complex thought process like human programmers would do for data flow analysis. These results suggest that LLMs do not have the same reasoning skill as human programmers.
@@
4.3 RQ3: Open Models, Static Analysis Tasks
@@
Table 8 showed results for AST generation in Java and C/C++. In Java dataset, we observed that open-source models jam reached Levenshtein score of 0.89 and CodeLlaMA reached Levenshtein score of 0.77 for AST generation. This result shows that LLMs learn meaningful information for AST generation during finetuning. We found that the result of CodeLlaMA is slightly lower. The possible explanation is that we used less data to finetune the model due to the computational cost although CodeLlaMA is already a larger model with more pretrained data compared with jam. To our surprise, we found that jam even outperforms gpt-in-context in this task although we cannot strongly claim that smaller models can do better than gpt-in-context on this task because we did not try to find the best prompt for the task. However, we found that an appropriate finetune helps to improve this task. For C/C++ we also observed the lower Levenshtein score compared with Java. This result aligns with the results from code development tasks that C/C++ has lower performance compared with Java in Section 4.1. Interestingly, we observed it is possible that LLMs make a very simple syntactic error. For example, we observed in Figure 2(b) in Area 1 that the AST generated by CodeLlaMA misses one “)” compared with Area 1 in Figure 2(c). Moreover, we observed in Figure 2(b) in Area 2 that LLMs might continue generating the token at the end. Overall, this shows the lack of structural information in the LLMs.
@@
We showed the result for both C/C++ and Java for callgraph generation in Table 9. We found that CodeLlaMA has high Jaccard similarity score and pair accuracy for callgraph generation after finetuning. However, we found that the chain accuracy is still very low. This result shows that models can only partially generate the correct callgraph. We also found that C/C++ dataset has lower score compared with Java dataset. The possible explanation is that we input the entire file for C/C++ dataset and this makes the problem harder. In addition, we only use 10k tokens during prediction due to hardware limitation. It is also possible that using entire files would require more tokens. However, the overall conclusion that lower chain accuracy and higher pair accuracy and Jaccrad similarity would remain the same based on our findings from Java. Overall, we found that LLMs struggle with looking at the longer context although the input size is long (Li et al., 2024a).
@@
The results for both C/++ and Java for data flow graph generation are in Table 9. Overall, we observed that the results for data flow graph generation has the similar results as in callgraph generation i.e. low chain accuracy but higher Jaccard similarity and pair accuracy and C/C++ dataset has lower results compared with Java. In Java, we also found that jam has very close results with CodeLlaMA on pair accuracy with the same finetuning data on data flow graph generation in Java dataset. This is to our surprise because larger models usually have bigger improvement in code development tasks e.g. 25% improvement on meteor score for code summarization in Java. However, it only shows a slight improvement or even worse on static analysis tasks. This result implies that for static analysis tasks, LLMs tend to saturate faster than other code development tasks. A possible explanation is that it is harder for LLMs to learn the useful information in static analysis tasks.
@@
To sum up, we found the similar results on both open-source models and closed-source models. Specifically, we found that LLMs have larger improvement on AST generation after finetuning. However, LLMs struggle on callgraph generation and data flow graph generation. This result shows LLMs’ inability on more complex reasoning tasks such as static analysis tasks.
@@
jam-base boolean function( double arg0){ return (( arg0 >= this.x ) && ( arg0 <= this.x + this.z) && ( ( arg0 >= this.y ) && ( arg0 <= this.
jam-srcml boolean function(double arg0, double arg1 ){ return (arg0 >= x) && (arg0 <= arg1); }
jam-dataflow boolean function(Coordinate arg0 ){ return ( x == arg0.x && y == arg0.y && z == arg0.z ); }
Table 13: Code generation example for C/C++.
Method ID 10
int binaryDec(int arr [], int n){
int ans = 0;
for(int i = 0 ; i < n ; i ++ )
ans += arr [ i ] << ( n - i - 1 );
return ans;
codellama-base int toDecimal(int arr[] , int n) { int res = 0 ; for(int i = 0; i < n;i ++){ res += arr [i] * ( 1 << i ); } return res; }
codellama-srcml int toDecimal(int arr [], int n){ int res = 0; for(int i = 0; i < n; i ++ ) res += arr [ i ] * ( 1 << i ); return res; }
codellama-dataflow int toDecimal(int arr [], int n){ int res = 0; for(int i = 0; i < n; i ++){ res += arr [i] * ( 1 << i ); } return res; }
codellama-callgraph int convertFromBinary(int arr []){ int n = 0 , i = 0 ; while(arr[i] != 0 ){ n = n * 2 + arr [i]; i ++; } return n; }
Table 14: Code generation example for C/C++.
Method ID 10
int EnneacontahexagonNum(int n){
return (94 * n * n - 92 * n) / 2;
codellama-base int enneacontahexagonNum(int n){ return ( 118 * n * n - 106 * n ) / 2; }
codellama-srcml int enneacontahexagonNum(int n){ return ( 11 * n * n - 10 * n ) / 2; }
codellama-dataflow int enneacontahexagonNum(int n){ return (11 * n * n - 10 * n) / 2; }
codellama-callgraph int enneacontahexagonNum(int n){ return (14 * n * n - 14 * n + 1) / 2; }
4.4 RQ4: Open Models, Static Analysis + Dev.
We showed the results for code summarization in Table 4 in Java and C/C++. We do not strongly find the statistical difference between the models pretrained with static analysis and the baseline model for both jam and CodeLlaMA for Java and CodeLlaMA for C/C++. The possible explanation is that the model looks at the keyword in the method for source code summarization instead of using static analysis tasks like what human programmers would do to comprehend source code. We showed an example in Java in Table 10. We found that LLMs tend to copy the technical jargon from the method for code summary. Specifically, we found that the summary from CodeLlaMA use “CoordinateAxis” in the summary even after static analysis tasks. Similarly, we also observed the same phenomenon in C/C++. For example, in Table 11, we found that human reference uses “render” as a verb to describe the method. However, the summary from CodeLlaMA uses “show” as a verb, which matches the signature of the function. Interestingly, we found that some of the summaries are the same. This is because we use very small temperature i.e. 1e-9 and models may generate the same result when the weight are similar (Peeperkorn et al., 2024), which further showed that pretraining models with static analysis tasks does not improve code summarization tasks significantly.
Table 5 shows the result for code generation tasks for both Java and C/C++. Overall, we do not strongly observe any statistical difference in Java and C/C++ dataset, i.e. the statistical significant difference across all static analysis tasks and two different programming languages. The possible explanation is that LLMs do not use static analysis tasks as the thought process. For example, we found in Table 4.3 that code generated by jam-srcml has better syntactic structure compared with jam-base. However, the underlying logic is still different from the reference. Another example in C/C++ dataset in Table 4.3 is that we observed that the code generated by CodeLlaMA without pretraining with any static analysis tasks has the logical error in the “toDecimal” method. After pretraining with static analysis tasks, we still found the same logical error. Interestingly, we observed that LLMs might generate a code with incorrect details. In Table 4.3, we found that CodeLlaMA does not generate the correct parameters for Enneacontahexagon numbers. For human being to do this task, they would look at the correct parameters before writing code. This implies that LLMs use a random number or a specific number from other information for some tasks. This shows the lack of reasoning skills in LLMs and LLMs do not use static analysis tasks for code generation.
We showed the results on code translation for both Java and C/C++ in Table 6. In Java, we do not observe statistical difference between the model pretrained with data flow analysis and srcml for both jam and CodeLlaMA although we do observe callgraph helps to improve the result on this task. However, we do not observe any statistical difference between in C/C++ dataset although we do observe the improvement when we pretrained the model with callgraph and data flow graph. The possible explanation is that pretraining with callgraph or data flow helps models look at the larger context and code translation would require a larger context for more accurate translation. However, we cannot strongly attribute the improvement to this because models only have partial improvement on callgraph and data flow generation. Another finding is that code translation have higher performance compared with another task in Java dataset. It is also possible that Java and C# have very similar API calls and syntax, so LLMs can easily copy the information from the source language as we discussed in RQ1 (refer to Section 4.1) as we do not observe similar trend in C to Rust dataset.
Overall, we do not strongly find that static analysis tasks lead to better results on coding tasks for both C/C++ and Java although we do find statistical significance in some cases. These lead to two conclusions 1) LLMs might have different thought process for coding tasks compared with human programmers 2) the high performance might suggest that LLMs use semantics information instead of syntactic information or more complex reasoning.
5 Conclusion
In this paper, we have three key observations. First, for code intelligence tasks, we observed that LLMs usually have higher performance on more popular programming languages. Specifically, we observed models usually have higher performance on Java compared with C/C++. The gap between these two programming languages may be because that Java has more training data (López et al., 2024). Moreover, we also found that the performance of LLMs is usually lower when the underlying structure of source code is more complicated. This shows that LLMs’ reasoning skill is limited.
Second, we found that in-context learning and finetuning only improve the performance on AST generation. However, in-context learning does not improve more complex static analysis tasks and finetuning only improves the more complex static analysis tasks partially. Specifically, we do not observe any significant improvement on callgraph generation and data flow graph generation with in-context learning. Moreover, we found that finetuning improves pair accuracy and Jaccard similarity although we do not observe the significant improvement on chain accuracy. This result shows that LLMs struggle with more complex reasoning.
Finally, we found that pretraining models with various static analysis tasks does not necessarily improve the code intelligence tasks. Specifically, we do not strongly observe the statistical difference across all datasets and all programming languages when we pretraining models with different static analysis tasks. Another evidence that showed that static analysis tasks do not improve code intelligence tasks is that LLMs tend to make the same logical errors after pretraining models with static analysis tasks. To sum up, this shows that LLMs have different thought process compared with human programmers.
Acknowledgements.
This work is supported in part by the NSF grants CCF-2100035, CCF-2211428, and CCF-2211429. Any opinions, findings, and conclusions expressed herein are the authors’ and do not necessarily reflect those of the sponsors.
Ethical approval
This study was conducted in accordance with institutional guidelines and approved by the Institutional Review Board.
Informed consent
Not applicable.
Author Contributions
Chia-Yi Su designed the experiments, implemented the code for the experiments, and wrote the manuscript. Collin McMillan wrote the manuscript and designed the experiments.
Code / Data Availability Statement
We release all data, scripts, survey materials, and results at https://github.com/apcl-research/llm-reason.
Clinical trial number
Not applicable
Conflict of interest
The authors declare that they have no conflict of interest.

### v2

@@ 3.9 Research Methods – RQ3
Our research method for answering RQ3 is to finetune the open-source LLMs on downstream tasks after those models have been finetuned on static analysis tasks. We show those configurations in Tables 1 and 2. For example, in the first row of jam column in RQ3, we show that we first finetune jam on the AST generation task. Then, we further finetuned it for source code summarization. We compared this model with the one without finetuning with static analysis tasks to show whether static analysis improves the performance of LLMs on code intelligence tasks.
The first row of Table 1 show that the dataset we use for code summarization in Java is 170k dataset in su2024distilled. However, Table 1 also shows that we did not use jam for C/C++ datasets because jam is pretrained for Java only (see Section 3.4). Our guiding principle in all fine-tuning and evaluation is the same as RQ2. We used the base jam and CodeLlaMa models released by their respective papers where those papers pretrained using large datasets of code (see Section 3.4). Then, we further trained those models with the training sets which were released with each dataset (see Sections 3.2 and 3.3). We use the hyperparameters as described in RQ2.
We evaluated each configuration using the metrics and test datasets in the papers for each dataset. For code summarization, we used the metrics METEOR and USE as implemented by su2024distilled and recommended by haque2022semantic. For code translation and code generation, we used Bleu score (papineni2002bleu) as in the dataset that we used for experiments (iyer2018mapping; lu2021codexglue). We also have additional datasets for code generation and translation evaluated with pass@k chen2021evaluating and compilation success rate.
@@ 3.10 Research Methods – RQ4
Our research method for answering RQ4 is to finetune models with code intelligence tasks and evaluate the models with static anlysis tasks. To this end, we used CodeLlama fintuned with each code intelligence to generate the three different static analysis tasks. We show the configuration in Table 2. For example, in the first row of RQ4, we show that we finetuned CodeLlama with code summarization. Then, we used the same model to generate AST. Note that we used simple prompts because our goal is to test the model’s capability instead of finding the best results with better prompts.
3.11 Research Methods – RQ5
Our research method for answering RQ5 is to classify the types of error for static analysis tasks based on the current literature. To this end, we classify the error types for callgraph into four categories, which are missing direct call, extra direct calls, missing indirect calls, reference cycle, and predicted cycle based on lehmann2023that; sui2020recalllehmann2023that constructed the callgraph that could potentially include cycles, direct edges, and indirect edges. We further classify into extra/missing direct calls and missing indirect calls based on false positive and false negative defined by murphy1998empirical, where false positive means additional edges and false negative means missing edges. For dataflow analysis, we followed weideman2025data to categorize the errors into direct edges and indirect edges. Based on false positive and false negative, we further classified these two edges into extra direct edges, extra indirect edges, missing direct edges, and missing indirect edges. We categorized the AST errors into missing nodes and extra nodes based on AST difference discussed by falleri2014finegrained; fan2021differential. In addition, alikhanifard2025novel showed that AST can be incorrect semantically, so we added tag mismatch into our categories. We also added parsing error because a generated AST may have an incorrect format/structure that causes the compile errors.
@@ 3.12 Threats to Validity
We organize threats to validity into threats to internal validity, external validity, and construct validity based on wohlin2012experimentation. The threats to internal validity are that different hyperparamters can have different results. We mitigate this threat by following the accompanying papers. The threats to construct validity are that our experimental results may not be able to generalize to different tasks and programming languages. We mitigate this by using three different code development tasks and static analysis tasks along with two different programming languages although more programming languages can have better generalizability of the findings. We leave the experiments for other programming languages as future work. The threats to external validity are that different models might generate different results. We mitigate this threat by using two different commercial LLMs and two different open-source LLMs.
3.13 Limitations
Our prompt design is the key limitation in this paper. We used the simple prompt to test the reasoning capabilities of LLMs. The results show that LLMs do not do well on static analysis tasks. However, these results may not be able to generalize to more advanced prompting strategies (e.g., agentic development) and reasoning models, such as GPT-o series. Note that we did not have access to GPT-o series models at the time we conducted this experiment although xie2025core show that more advanced models still struggle with tasks that require deeper semantic understanding.
@@ 4 Experimental Results
dataflow graph gpt-base 0.0047 0.0111 0.0
gpt-in-context 0.0494 0.0951 5.37E-03
gpt-in-context-reason 0.0479 0.0941 4.53E-03
gemini-base 0.0045 0.0114 0.0
gemini-in-context 0.0522 0.1028 2.31E-02
gemini-in-context-reason 0.0711 0.1426 3.58E-03
codellama-finetuned 0.54 0.74 2.03E-01
jam-finetuned 0.38 0.54 0.18
@@
dataflow graph gpt-base 0.0017 0.0054 1.12E-04
gpt-in-context 0.0059 0.0126 5.13E-03
gpt-in-context-reason 0.0047 0.0118 3.91E-03
gemini-base 0.0022 0.0056 1.59E-03
gemini-in-context 0.0093 0.0176 9.62E-03
gemini-in-context-reason 0.0047 0.0118 3.91E-03
codellama-finetuned 0.12 0.25 1.02E-01
@@
4.1 RQ1: Closed Models, Static Analysis Tasks
We found that in-context learning improves AST generation tasks and popular programming languages have higher improvement. We showed the results for AST generation on both Java and C/C++ in Table 4. In Java, we found that gpt-base has the Levenshtein score of 0.45 and gemini-base has the Levenshtein score of 0.8. We also found the similar results in gemini although we found higher results in gemini, which has Levenshtein score of 0.64 for gemini-base and Levenshtein score of 0.93 for gemini-in-context. For C/C++, we observed the similar trend as in Java. However, we observed the smaller improvement in C/C++ dataset. For example, we found around 80% improvement between GPT-base and GPT-in-context in Java dataset. However, we only found around 30% improvement between GPT-base and GPT-in-context in C/C++ dataset, which has around 50% difference in improvement. This result aligns with the finding that LLMs can have better performance on more popular programming languages (jimenez2024evaluation). The gap between C/C++ and Java may imply the lack of reasoning skills on LLMs to generalize from one programming language to another programming language.
@@ 4.1 RQ1: Closed Models, Static Analysis Tasks
We found that LLMs do a poor job on callgraph and dataflow graph generation and in-context learning does not improve the results. We showed the results for callgraph and dataflow graph in Table 5 for both Java and C/C++. For callgraph generation in Java, we found that gpt-base has Jaccard similarity of 0.121 and gpt-in-context has Jaccard similarity of 0.122. gemini-base and gemini-in-context have slightly higher performance, which is 0.186 and 0.191 respectively. We also found that chain accuracy is the lowest among all other metrics. In C/C++, we found the similar trend that in-context learning does not help to improve the results and chain accuracy is the lowest among all other metrics. For data flow generation, we found the similar results on both Java and C/C++ dataset. This result indicates that LLMs do not do well on callgraph and dataflow graph generation.
We found that reasoning prompts do not improve the performance for both callgraph and dataflow graph generation. We show the results for reasoning prompts in Table 5. For example, we observe in Java that gpt-in-context has Jaccard similarity of 0.122, but we only observe 0.113 in gpt-in-context-rea- son. The possible explanations are 1) LLMs do not have the fundamental ability for callgraph and dataflow generation, 2) in-context learning is only helping models to learn the format instead of the knowledge to do a task (long2024does), and 3) models still struggle to comprehend longer prompt although LLMs can take very long context (zamfirescu2023why; liu2024lost). A possible improvement would be guiding LLMs with more detailed instructions on each task instead of providing examples.
Overall, we found that closed-source LLMs do not do static analysis very well. To our surprise, we found that in-context learning only helps to improve AST generation but not callgraph generation and dataflow generation. The possible explanation is that it is easy for LLMs to learn the tags for AST generation. However, LLMs struggle with looking at a larger context for callgraph generation and more complicated reasoning tasks. These results suggest that LLMs do not have the same reasoning skills as human programmers.
@@
4.2 RQ2: Open Models, Static Analysis Tasks
@@
We observed that LLMs learn meaningful information after finetuning. Table 4 shows the results for AST generation in Java and C/C++. In Java dataset, we observed that open-source models jam reached Levenshtein score of 0.89 and CodeLlaMA reached Levenshtein score of 0.77 for AST generation. We also found that the result of CodeLlaMA is slightly lower. The possible explanation is that we used less data to finetune the model due to the computational cost although CodeLlaMA is already a larger model with more pretrained data compared with jam. To our surprise, we found that jam even outperforms gpt-in-context in this task in the simple prompt settings. This shows that an appropriate finetuning helps to improve this task. For C/C++, we also observed the lower Levenshtein score compared with Java. Interestingly, we observed that LLMs can make a very simple syntactic error. For example, we observed in Figure 2(b) in Area 1 that the AST generated by CodeLlaMA misses one “)” compared with Area 1 in Figure 2(c). Moreover, we observed in Figure 2(b) in Area 2 that LLMs might continue generating the token at the end. This shows the lack of semantic understanding in LLMs.
@@
We observed that models can only partially generate the correct callgraph. We showed the result for both C/C++ and Java for callgraph generation in Table 5. We found that CodeLlaMA has high Jaccard similarity score and pair accuracy for callgraph generation after finetuning. However, we found that the chain accuracy is still very low. We also found that C/C++ dataset has lower score compared with Java dataset. The possible explanation is that we input the entire file for C/C++ dataset and this makes the problem harder. In addition, we only use 10k tokens during prediction due to hardware limitation. It is also possible that using entire files would require more tokens. However, the conclusion is that lower chain accuracy and higher pair accuracy and Jaccrad similarity would remain the same based on our findings from Java. This indicates that LLMs struggle to look at the longer context although the input size is long (li2024long).
@@
We observed that the results of dataflow generation are similar to callgraph generation, i.e., low chain accuracy but higher Jaccard similarity and pair accuracy and C/C++ dataset has lower results compared with Java. The results for both C/C++ and Java for dataflow graph generation are in Table 5. In Java, we also found that jam has very close results with CodeLlaMA on pair accuracy with the same finetuning data on data flow graph generation in Java dataset. This is to our surprise because larger models usually have bigger improvement in code development tasks, e.g., 25% improvement on meteor score for code summarization in Java. However, it only shows a slight improvement or even worse on static analysis tasks. This result implies that LLMs saturate faster in static analysis tasks compared with other code intelligence tasks. A possible explanation is that it is harder for LLMs to learn the meaningful information in static analysis tasks.
@@
To sum up, we found the similar results on both open-source models and closed-source models. Specifically, we found that LLMs have larger improvement on AST generation after finetuning. However, LLMs struggle with callgraph generation and dataflow graph generation. This result shows LLMs’ inability on static analysis tasks.
@@
jam-base boolean function( double arg0){ return $($ arg0 >= this.x ) && ( arg0 <= this.x + this.z) && ( ( arg0 >= this.y ) && ( arg0 <= this. jam-srcml boolean function(double arg0, double arg1 ){ return (arg0 >= x) && (arg0 <= arg1); }
jam-dataflow